OpenStudy (paigeweeks):

Les bought 6 pairs of shorts for s dollars each and a blazer that costs three times as much as a pair of shorts. He spent a total of $139.50. How much did the blazer cost ? Equation :____________ Blazer cost:________________

OpenStudy (markchernioglo):

Your equation if you think about it is 6s+18=139.50.

OpenStudy (markchernioglo):

Someone correct me if im wrong please.

OpenStudy (markchernioglo):

Now to find the cost we do the math... We gotta find s by itself. So what we do is subtract 18 from 139.50 Will look like this: 6s=121.5 Then finally we divide 6 by 121.5 so 121.5/6=20.25 So s=20.25

OpenStudy (paigeweeks):

Then, I have to multiply by 3 right?

OpenStudy (markchernioglo):

No I already put that into the equation I did 6*3 and got 18.

OpenStudy (paigeweeks):

Oh okay ! Thanks

OpenStudy (markchernioglo):

Np Paige good luck! :)

OpenStudy (paigeweeks):

Thanks

OpenStudy (sushi121212):

No wait, isn't it 6s+18s=139.50 because the blazer costs 3 times as much as the shorts?

OpenStudy (markchernioglo):

But then think about it there is no point of adding an s it will still be distributed. There will be one s again.

OpenStudy (sushi121212):

Am I wrong? If s = the money, and 6 pair of pants is 6s, then the blazer that is 3 times the amount of the pair of pants would be 18s. Yes, the s will add up to 24s=139.50 as the equation

OpenStudy (sushi121212):

If the equation is 24s=139.50, then the amount for a blazer should be 5.8125.

OpenStudy (markchernioglo):

Yea, I screwed up I guess your right. I totally lost my mind there lol.

OpenStudy (markchernioglo):

But wait... we both are probably wrong because it does NOT say the cost of the shorts...

OpenStudy (markchernioglo):

@Sushi121212

OpenStudy (markchernioglo):

6s+1s*3=139.50

OpenStudy (sushi121212):

Well, if you set it up like this: 24s=139.50 Then s=5.8125 Then you can find the cost of the shorts: 6(5.8125) = 34.88 Then multiply it by 3: 3 (34.88) = 104.64 I don't know how to do it on one equation though.

OpenStudy (markchernioglo):

My question is how did you get 24?

OpenStudy (markchernioglo):

Because remember we don't have the value of s in the first place, thus we can't say 6s+18s=139.50

OpenStudy (sushi121212):

We're trying to find the value of s so we can use it to find the price of the blazer.

OpenStudy (markchernioglo):

So how can we say 18?

OpenStudy (sushi121212):

Wait, but: 6s+[6s(3)]=139.50 Is the same thing as: 6s+18s=139.50 24s=139.50

OpenStudy (markchernioglo):

Look at the question carefully, it doesn't say 6 blazers. Its says: "and a blazer that costs three times as much as a pair of shorts."

OpenStudy (sushi121212):

That's why I put it as 6s(3)

OpenStudy (markchernioglo):

That is why you can't use 6 lol.

OpenStudy (markchernioglo):

It says a blazer meaning one. :D

OpenStudy (markchernioglo):

and its not saying 3 times the amount, but the cost.

OpenStudy (markchernioglo):

Sorry Paige this got way more confusing then how it looks lol.

OpenStudy (sushi121212):

Oh, then it's 3s? So it would be 6s+3s = 139.50? Or 9s = 139.50?

OpenStudy (sushi121212):

s is the price of one pair of shorts, right?

OpenStudy (markchernioglo):

Yea!

OpenStudy (markchernioglo):

That makes sense now.

OpenStudy (markchernioglo):

So would probably be 3s

OpenStudy (sushi121212):

FInally!

OpenStudy (markchernioglo):

Yea! Ikr? haha

OpenStudy (sushi121212):

Well, at least we solved it in the end! I hope PaigeWeeks didn't write the wrong answer...

OpenStudy (markchernioglo):

Lol yea me too hope the best for Paige.